Goshford Company produces a single product and has capacity to produce 100,000 units per month. Costs to produce its current sales of 80,000 units follow. The regular selling price of the product is $146 per unit. Management is approached by a new customer who wants to purchase 20,000 units of the product for $77.40 per unit. If the order is accepted, there will be no additional fixed manufacturing overhead and no additional fixed selling and administrative expenses. The customer is not in the company’s regular selling territory, so there will be a $7.20 per unit shipping expense in addition to the regular variable selling and administrative expenses. Per Unit Costs at 80,000 Units Direct materials $ 12.50 $ 1,000,000 Direct labor 15.00 1,200,000 Variable manufacturing overhead 10.00 800,000 Fixed manufacturing overhead 17.50 1,400,000 Variable selling and administrative expenses 16.00 1,280,000 Fixed selling and administrative expenses 15.00 1,200,000 Totals $ 86.00 $ 6,880,000 Calculate the combined total net income if the company accepts the offer to sell additional units at the reduced price of $77.40 per unit. Loading... Determine whether management should accept or reject the new business. Accept Reject
Answer #1

Calculate combined Total Net income :

80000 Units 20000 Units
Sales 80000*146 = 11680000 20000*77.4= 1548000
Cost
Direct material 1000000 20000*12.5 = 250000
Direct labor 1200000 20000*15 = 300000
Variable manufacturing overhead 800000 20000*10 = 200000
Fixed manufacturing overhead 1400000 0
Variable selling and administrative expense 1280000 23.2*20000 = 464000
Fixed selling and administrative expense 1200000 0
Total Cost 6880000 1214000
Net operating income 4800000 334000

Combined total net operating income = 4800000+334000 = $5134000

Yes company should accept the new bussiness